OpenStudy (anonymous):

What is the range of:

OpenStudy (anonymous):

\[f ^{-1}\] when \[f(x)=\sqrt{x-5}\]

OpenStudy (anonymous):

And first, find \[f ^{-1}\]

OpenStudy (solomonzelman):

range of the inverse or of f(x) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But first find the inverse of f(x)

OpenStudy (akashdeepdeb):

y = \(\sqrt{x-5}\) \(y^2 = x - 5\) x = \(y^2 + 5\) \(f^{-1} (x) = x^2 + 5\) Domain of f equals range of inverse of f. Find the domain of f.

OpenStudy (anonymous):

I don't know how do do that part

OpenStudy (solomonzelman):

I will draw it...

OpenStudy (anonymous):

domain of f(x) = All real numbers 4<?

OpenStudy (solomonzelman):